lim x→0 (e tanx - e x ) / (tan x -x) is equal to? (A) 1 (B) 1/2 (C) 1/3 (D) 0 [BOOK ANSWER: 1/2] EXPLAIN .

lim x→0   (etanx -  ex) / (tan x -x) is equal to?


(A) 1 (B) 1/2 (C) 1/3 (D) 0


[BOOK ANSWER: 1/2]


EXPLAIN .

Grade:12th Pass

3 Answers

Aman Bansal
12 years ago

Dear Aditi,

Expand tan x using taylors theorem and then reduce the expression to get the answer.

Best Of Luck...!!!!

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

Chetan Mandayam Nayakar
12 years ago

the Mcclaurin series expansion is

ey=∑y=0 to ∞yn/n!

therefore, lim x→0   (etanx -  ex) / (tan x -x)=1+(1/(tanx-x))(tan2x-x2+tan3x-x3...

=1+(1/(tanx-x))(tan2x/(1-tanx))-(x2/(1-x))=1+tanx+x=1

In case more explanation is needed, please email me at:

 

jitender lakhanpal
12 years ago

Dear Aditi

 

lim x→0   (etanx -  ex) / (tan x -x)

take ex   as common

so we get

lim x→0 ex  (etanx-x -  1) / (tan x -x)

this is nothing but the standard limit

limit x tends to 0 (ex - 1)/x = 1 

where x = tanx - x 

so it is equal to 1

so the answer is 1 and not 1/2   

we can also check this by L hospital rule as it is 0/0

 

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

jitender